OpenStudy (anonymous):

Solve the equation 2/3 (x − 2) = 4x for x.

OpenStudy (yrelhan4):

[2(x-2)]/3=4x.. is this the question?

OpenStudy (anonymous):

2 over 3 (x-2) = 4x for x

OpenStudy (yrelhan4):

x-2 in numerator or denominator?

OpenStudy (anonymous):

well look maybe this might help you have to find x Solve the equation 2 over 3 (x − 2) = 4x possible answers 2 over 5 negative 5 over 2 negative 2 over 5 positive 5

OpenStudy (aravindg):

Wait are you taking a test ?

OpenStudy (anonymous):

no its a practice for test i need to understand how to do it and then it will help me for my real test

OpenStudy (anonymous):

so can you help me ?

OpenStudy (kenljw):

[2(x-2)]/3=4x 2x-4=12x 10x=4 x=2/5

OpenStudy (aravindg):

ok First and foremost Multiply by 3 on both sides

OpenStudy (anonymous):

yes

OpenStudy (aravindg):

Did you understand how to get the answer?

OpenStudy (anonymous):

x= 2/5 ?

OpenStudy (aravindg):

I got -2/5

OpenStudy (aravindg):

2x-4=12x 10x=-4 x=-2/5

OpenStudy (kenljw):

yes my error

OpenStudy (anonymous):

i have recheck it -2/5

OpenStudy (anonymous):

thank you for your help both of you thanks